# Multi-User Chat Assistant (MUCA): a Framework Using LLMs to Facilitate Group Conversations

### 10 Jan 2024 v1

@@ 4 Evaluation
We conducted case studies and user studies to examine the MUCA’s performance in group conversations with various discussion topics and group sizes. Case studies describe the objective strengths of our design, while the user studies show the subjective advantages of the MUCA which can be easily identified by users. As to MUCA implementation, we refer the readers to Appendix A.2.5 for implementation details.
@@ 4.1.1 MUCA Configurations
In this section, we evaluated our proposed MUCA with slightly different configurations for various group sizes. A general description of the three MUCAs  [^5] In any experiments and results in this section, aliases _bot_Spirit, _bot_Perseverance, and _bot_Discovery were given to MUCA-basic, MUCA-advanced-small, and MUCA-advanced-medium, respectively. This renaming strategy is used to ensure that participants in user studies do not have prior knowledge of each chatbot, thereby preventing biases. is as follows:
MUCA-basic: we directly transfer single-user chatbot and its conceptual behavior to a multi-user setting, such baseline system we called MUCA-basic. It is implemented by GPT-4 with a single prompt, which takes user-input information, conversation context, and participants’ names as input and outputs generated responses. In the prompt, we simply define its dialog acts, for example, keep silent, direct chatting, and in-context chime-in. The MUCA-basic is applied and evaluated in a 4-person group chat with an execution interval ($N_{exe}$) of 3 and a short-term context window size ($N_{sw}$) of 8.
MUCA-advanced-small: it is applied in a 4-person group chat, and uses the same configuration ($N_{exe},N_{sw}$), and the same user-input information as MUCA-basic. The details of MUCA framework and functionalities are described in Section 3.
MUCA-advanced-medium: it is applied in an 8-person group chat and shares the same framework and architecture as MUCA-advanced-small but has different configurations. These configurations are automatically adjusted based on the number of participants ($N_{exe}=0.75*N_{p}$, $N_{sw}=2*N_{p}$) to maintain the latency-efficiency.
Figure 6: Illustration of conversational scenarios involving MUCA-advanced-small as a good assistant. It showcases the chatbot’s functionalities of chiming in at the right timing with appropriate content, along with its capabilities to accurately identify and respond to the relevant user. The sub-figures show the chatbot’s interactions: A) Direct chatting with the user who pinged it; B) Providing insights for decision making; C) Providing insights to advance a stuck conversation; D) Answering a general question; E) Encouraging lurkers’ participation; F) Summarizing past information automatically; G) Transiting the sub-topic when it’s 1) well-discussed or 2) uninteresting, respectively.
@@ 4.1.2 Dialog Topics
While the proposed MUCA has been designed to be a generic framework, for this evaluation, we focus on 4 goal-oriented communication tasks (i.e., estimation, decision making, problem-solving, and open discussion), rather than chit-chat or joke-making, which will be shown in Appendix A.2.6. The importance of a multi-user chatbot can be better reflected in goal-oriented communication tasks with our designed evaluation benchmarks.
@@ 4.2 Case Study
We qualitatively demonstrate our general observations for the proposed MUCA-advanced-small within the following use case studies.
Figure 7: Qualitative comparison between MUCA-basic and MUCA-advanced-small through case studies. A) and B) Comparisons of hallucination issues. C) Comparison of summarization feature. D) Comparison of conflict resolution capability.
@@ 4.2.1 Functionality Demonstration
In Figure 6, we showcase the main functionalities of our proposed MUCA-advanced-small to facilitate multi-user communication.
@@
Direct Chatting: In Figure 6-A), the interaction between Amy and MUCA-advanced-small is depicted where Amy used keyword @mubot to trigger the MUCA to respond immediately (as described in Section 3.2.3). Upon addressing Amy’s concern about the insurance fee, MUCA’s feedback not only validated her perspective but also offered supplementary insights to elucidate and support her argument. This clarification enhanced Amy’s viewpoint, thereby promoting a more comprehensive grasp of her statement among the other participants in the discussion.
@@
In-Context Chime-in (provide insights for decision-making): In Figure 6-B), it is important to observe that although MUCA-advanced-small was not directly pinged by any participant, it could perceive that the discussion involved selecting between a public park and a learning center for hosting an event. Also, it highlighted other important factors that were not explicitly mentioned by the participants. By doing so, it offered insights that encouraged deeper and more comprehensive consideration in the decision-making process.
@@
In-Context Chime-in (provide insights for a stuck conversation): In Figure 6-C), all the participants were indecisive and unsure about the location of the event, leading to a situation where the conversation was essentially stuck, as indicated by the backend log of MUCA-advanced-small. Once recognizing this, MUCA provided pros and cons for two venues and tried to advance the conversation and facilitate its forward movement.
@@
In-Context Chime-in (answer general questions): In Figure 6-D), Cindy raised a general question which is not for specific participants. By recognizing this, MUCA-advanced-small proactively chimed in to provide an answer, which helped the conversation move on to the next phase. In contrast to a single-user system where a chatbot invariably responds to a user’s query, a multi-user chatbot faces a more complex challenge: it has to discern whether to reply to questions when not directly pinged. This necessitates a more subtle decision-making process for the chatbot in a multi-user environment, as described in Section 3.1.2.
@@
Participation Encouragement: In Figure 6-E), it is observable that while Amy, Dennis, and Bob were actively engaged in the discussion, Cindy remained silent for an extended period of time. MUCA-advanced-small identified the inactive status of Cindy and hence decided to encourage her participation based on Cindy’s chatting record. This dialog act tries to improve Cindy’s engagement and to collect diverse opinions for a more informed decision on the topic. We emphasize here that the MUCA, by our design, is programmed to avoid excessively pinging Cindy or any user who appears reluctant to share their ideas, thereby ensuring a balanced and non-intrusive interaction, as described in Section 3.2.3.
@@
Initiative Summarization: Initiative summarization is generated automatically by the MUCA after 12 turns to conclude the existing conversation. In Figure 6-F), we can observe that MUCA-advanced-small triggered this function and compiled key points from chatting history. In this example, MUCA not only summarized the diverse opinions of each participant but also offered its own viewpoint, trying to help participants reach an agreement. It empowers participants to quickly catch up on the long and ongoing discussion so that they can contribute their thoughts without the need to review extensive chatting history.
@@
Sub-topic Transition: In Figure 6-G)-1 and G)-2, MUCA-advanced-small collected information during the chat, and attempted to shift the topic at an appropriate moment. As mentioned in Section 3.2.3, this dialog act would be triggered in two conditions: (1) As shown in Figure 6-G)-1, currently only a limited number of participants were engaged in a topic that had previously been extensively discussed by the majority, it would attempt to guide the conversation towards the next phase. (2) As shown in Figure 6-G)-2, if MUCA detected that the current topic had been discussed by only a small number of participants as Amy and Bob have not participated from the beginning, it then attempted to transit the topic to improve user engagement. In both scenarios, MUCA was customized to ping inactive participants and tried to boost their interest according to their chatting history before transiting the topic. Besides, the transition was not abrupt or forceful; it politely inquired whether participants wished to continue or shift the discussion, thus minimizing the chance of any potential disruption. Note that MUCA is designed to guide rather than dictate the flow of dialogue, ensuring that the conversation remains participant-driven.
@@ 4.2.2 Functionality Comparison
As shown in Figure 7, we compared the behaviors of proposed MUCA-advanced-small and those of the naive MUCA-basic to demonstrate the effectiveness of the design.
@@
Hallucination: As shown in Figure 7-A)-1 and A)-2, MUCA-basic made up some assumed information beyond the user-input information (topics, hints, and agenda), such as budget limit, participants’ age, and their genders to answer the questions. These assumptions may lead to user distrust, wrong decisions, and even bias. On the contrary, MUCA-advanced-small kindly notified the users that the question was out of scope, and its responses were aligned with the user input information.
@@
We dive deeper into this issue in Figure 7-B)-1 and B)-2. For the unknown budget information, MUCA-basic at first provided a non-existing number, which was later questioned and corrected by Bob who may know the real budget. Later, when Cindy tried to figure out who may know the budget, MUCA-basic still insisted on this hallucinated information, and even tried to transfer the topic. In contrast, for MUCA-advanced-small, it inferred that Bob may know the information, as he actively discussed this information. We are highlighting that, unlike single-user chatbot, multi-user chatbot has to process more complicated history, participants’ relationship, and user interactions, resulting in bigger challenges for the chatbot to extract useful information and generate hallucination-free response without dedicated prompting design, even using a powerful LLM as its backend model.
@@
Direct Chatting (passive summarization): Summarization or similar tasks (such as voting) are requested from users, especially when participants have different opinions in a long conversation. We define this request as passive summarization, in contrast with the Initiative Summarization as mentioned in Section 3.2.3. We compare this functionality in Figure 7-C)-1 and C)-2. For MUCA-basic, it failed to understand the query intent from Dennis, which was summarizing the votes from all participants. Instead, it summarized opinions, however, its summary was inaccurate due to the limited context window by design. For example, it mentioned the "Majority" leaning towards the learning center, where only Dennis voted for this option. On the contrary, MUCA-advanced-small, who enables to go beyond the limit of the window size, correctly summarized the vote and categorized them by different participants.
@@
Figure 8: A detailed comparison from the user study between the MUCA-basic and the MUCA-advanced-small. The left two and the right two columns represent two topics, indoor courses (Topic-A) and interview process (Topic-B), respectively. Each set of results presents the performance of MUCA-basic in the first row, and the performance of MUCA-advanced-small in the second row.
@@
Conflict Resolution: It is highly possible that participants have diverse opinions in a multi-user chatting environment, while it is not a very common scenario for a single-user setting, since the user generally has the same behavior pattern. As shown in Figure 7-D)-1, MUCA-basic tried to resolve the conflict by instructing users with its own opinion, which was very misleading and biased. Even worse, after giving its own opinion, MUCA-basic transferred to another topic, which may make users annoyed as it destroyed the conversation flow. In Figure 7-D)-2, MUCA-advanced-small wrapped up the conflict by different parties, and raised inspiring questions for people to think thoroughly, and resolved conflict when possible.
@@ 4.3 User Study
In this section, we conducted user studies to qualitatively and quantitatively demonstrate the effectiveness of the proposed MUCA-advanced-small and MUCA-advanced-medium, in comparison with MUCA-basic.
@@ 4.3.1 Study Design and Procedure
To investigate MUCA’s effectiveness, we conducted user studies with 3 groups of participants, i.e., two small groups and one medium group. Each small group consisted of 4 participants, and the medium one included 8 participants (females and males were 1:1 for each group). We chose two representative goal-oriented topics for the user studies, as shown in Section 4.1.2. The experiments on two small groups were used for comparison between the MUCA-basic and the proposed MUCA-advanced-small. The experiments on the medium group demonstrated the capabilities of MUCA on more complex chatting scenarios in a larger group.
For small-group experiments, the major task was to compare the proposed chatbot (MUCA-advanced-small) with the baseline chatbot (MUCA-basic). Group-A conducted experiments in the following order:
  • Topic A with 1) MUCA-basic and 2) with MUCA-advanced-small,
  • Topic B with 1) MUCA-advanced-small and 2) with MUCA-basic,
while Group-B conducted experiments by altering the order of the two chatbots for each topic. It was worth highlighting that the order in which the two chatbots were used would cause a learning effect. Participants may get more familiar with the topic after chatting with the first chatbot, resulting in smoother communication with the second chatbot. Therefore, when calculating the statistics for a specific topic, the reversed order in Group-A and Group-B would ensure that the order of chatbots did not bring bias.
We also applied the MUCA to a medium group using Topic-A to demonstrate the proposed chatbot (i.e., MUCA-advanced-medium) could be generally applicable to a larger group.
4.3.2 Qualitative Comparison with Small-size Groups
Statistics from users: We used the two small groups to compare the proposed MUCA-advanced-small and the baseline MUCA-basic. The detailed comparison can be found in Figure 8.
This figure consists of four sets of results (a) - (d). The first three sets are the statistics of responses to multi-choice questions, categorized according to the chatbots’ (a) chiming-in timing, (b) chiming-in content, and (c) participation encouragement. The set (d) is the overall evaluation scores.
The left two and the right two columns represent two topics, indoor courses (Topic-A) and interview process (Topic-B), respectively. Each set of results presents the performance of MUCA-basic in the first row, and the performance of MUCA-advanced-small in the second row. MUCA-advanced-small generally performances better compared with MUCA-basic.
@@ 4.3.2 Qualitative Comparison with Small-size Groups
Figure 9: General feedback from users for comparison between the MUCA-advanced-small and the MUCA-basic.
@@
Chime-in Timing: Participants have observed that both chatbots have ever chimed in at the right time during the whole conversation, while MUCA-advanced-small performs slightly better, as demonstrated in Figure 8(a). However, MUCA-basic also has been observed to frequently chime in at the wrong timing during the chat, unlike MUCA-advanced-small, which exhibits this behavior much less frequently. Note that 56.25% (9 out of 16) participants observed that MUCA-basic chimes in excessively. This is believed to be a result of its less strategically designed behavior – it always replies every three turns ($N_{exe}=3$) and ignores the "keeping silent" instruction in its prompt, as described in Section 4.1.1. In contrast, MUCA-advanced-small has not exhibited this issue as 0% (0 out of 16) participant report it. However, some participants observe that MUCA-advanced-small sometimes chime in insufficiently. This issue might stem from its pre-defined execution interval to maintain latency-efficiency as well as the dialog act of keeping silence, which causes it to reply every three turns at the most if not directly pinged by participants. Although the $N_{exe}$ could be reduced for MUCA-advanced-small, this would likely increase both response latency and resource usage, presenting a typical trade-off between resource and performance. We think we may also need more market investigation in the future to decide the best choice of $N_{exe}$ for a production-level design of a multi-user chatbot. One more reason for insufficient chime-in concluded through our observation is as below. In text-based communication systems, unlike speech-based systems, participants might overlook information provided by the chatbot, as evidenced in some chat logs. This oversight can lead to a gap between the information the chatbot has provided and the participants’ awareness. Therefore, sometimes users expect the chatbot to repeat information it has already given.
@@
Chime-in Content: Besides chiming in with good timing, MUCA-advanced-small usually produces appropriate content, as shown in Figure 8(b), and inappropriate content is not commonly observed. However, MUCA-basic repeats the information it mentions before, asks redundant questions, and tends to generate excessive content. Some of the provided information might be useful but some of them are overwhelmed, which might waste participants more effort to distinguish them.
@@
Participation Encouragement: The interaction feature, i.e., pinging a lurker by a chatbot, should be cautiously designed, including its timing, frequency, and contents, as it may impose negative feelings on some participants due to different personalities, while a good design may improve the efficiency of the chatting and make users more engaged. From this user-friendly perspective, as shown in Figure 8(c), MUCA-advanced-small generally has better behaviors with the proposed design, as it improves comfortableness and reduces users’ negative feelings.
@@
As we can see, MUCA-advanced-small shows notable successes in most cases as demonstrated in Figure 8(d). From the perspective of the three major user-friendly factors, MUCA-advanced-small received significantly higher scores.
Indoor Course (Topic-A) Interview (Topic-B)
M*-b M*-adv-sml M*-adv-med M*-b M*-adv-sml
Quantitative Results
Engagement (words ex.) 974 1239 1212 1186 1393
Engagement(Avg. msg len.) 8.18 8.91 10.54 9.64 10.47
Evenness 122 $\pm$ 70 155 $\pm$ 66 152 $\pm$ 69 148 $\pm$ 77 174 $\pm$ 96
Consensus 0.5 0.67 0.67 0.5 1.0
User Attitudes
Efficiency 3 3 4 0 4
Conciseness 1 3 2 -6 7
Usefulness 5 6 4 -2 7
Overall rate 4.375 7.25 7.25 5.75 6.75
Table 1: Comparisons among different chatbots with two topics. M*-b, M*-adv-sml and M*-adv-med represent MUCA-basic, MUCA-advanced-small and MUCA-advanced-medium, respectively.
A more straightforward comparison can be found in Figure 9. We can find that MUCA-advanced-small generally performs better from three evaluation perspectives as concluded from the figure. First, the MUCA-advanced-small can chime in at the right timing without disturbing the participants. Second, the proposed MUCA-advanced-small is a good assistant, which can summarize topics, help the user understand the topics, reply to users’ requests, and improve their engagements. Third, the MUCA-advanced-small is a good leader to lead and transfer topics.
4.3.3 Quantitative Study with Small-size Groups

### 16 Feb 2024 v3

@@ 4 Evaluation
We conducted case studies and user studies to examine the MUCA’s performance in group conversations with various discussion topics and group sizes. Case studies describe the objective strengths of our design, while the user studies show the subjective advantages of the MUCA which can be easily identified by users. As to MUCA implementation, we refer the readers to Appendix A.4 for implementation details.
@@ 4.1.1 MUCA Configurations
This section evaluates a baseline model based on GPT-4 OpenAI (2023) and our proposed MUCAs with slightly different configurations for various group sizes. A general description of the baseline system and two proposed MUCAs  [^5] In this section, aliases _bot_Spirit, _bot_Perseverance, and _bot_Discovery were given to Baseline-small, MUCA-small, and MUCA-medium, respectively. It ensures that participants in user studies do not have prior knowledge of each chatbot, thereby preventing biases. is as follows:
Baseline-small: We directly transfer single-user chatbot and its conceptual behavior to a multi-user setting, such baseline system we called Baseline-small. It is implemented by GPT-4 with a single prompt, which takes user-input information, conversation context, and participants’ names as input and outputs generated responses. In the prompt, we simply define its dialog acts, for example, keep silent, direct chatting, and in-context chime-in, as shown in Appendix  A.5. The Baseline-small is applied and evaluated in a 4-person group chat with an execution interval ($N_{exe}$) of 3 and a short-term context window size ($N_{sw}$) of 8.
MUCA-small: It is applied in a 4-person group chat, and uses the same configuration ($N_{exe},N_{sw}$), and the same user-input information as Baseline-small. The details of MUCA framework and functionalities are described in Section 3.
MUCA-medium: It is applied in an 8-person group chat and shares the same framework and architecture as MUCA-small but has different configurations. These configurations are automatically adjusted based on the number of participants ($N_{exe}=0.75*N_{p}$, $N_{sw}=2*N_{p}$) to maintain the latency-efficiency.
Figure 6: Illustration of conversational scenarios involving MUCA-small. It showcases the chatbot’s functionalities of chiming in at the right timing with appropriate content, along with its capabilities to accurately identify and respond to the relevant user. The sub-figures show the chatbot’s interactions: A) Direct chatting with the user who pinged it; B) Providing insights for decision-making; C) Providing insights to advance a stuck conversation; D) Answering a general question; E) Encouraging lurkers’ participation; F) Summarizing past information automatically; G) Transiting the sub-topic when it’s 1) well-discussed or 2) uninteresting, respectively. The conversation consists of 1 chatbot and 4 users, namely, Amy, Bob, Cindy, and Dennis.
@@ 4.1.2 Dialog Topics
While the proposed MUCA has been designed to be a generic framework, for this evaluation, we focus on 4 goal-oriented communication tasks (i.e., estimation, decision making, problem-solving, and open discussion), rather than chit-chat or joke-making, which will be shown in Appendix A.6. The importance of a multi-user chatbot can be better reflected in goal-oriented communication tasks with our designed evaluation benchmarks.
@@ 4.2 Case Study
We qualitatively show MUCA-small’s key functions using case studies, which are created by MUCA interacting with designed user inputs. We demonstrate our general observations for the MUCA-small within the following case studies.
Figure 7: Qualitative comparison between Baseline-small and MUCA-small through case studies. A) and B) Comparisons of hallucination issues. C) Comparison of summarization feature. D) Comparison of conflict resolution capability.
@@ 4.2.1 Functionality Demonstration
In Figure 6, we showcase the main functionalities of our proposed MUCA-small to facilitate multi-user communication introduced in Section 3.2, including direct chatting, in-context chime-in for helping decision, advancing stuck conversation, and answering general questions, participation encouragement, initiative summarization, and sub-topic transition.
@@
Direct Chatting: In Figure 6-A), the interaction between Amy and MUCA-small is depicted where Amy used the keyword of @mubot to trigger the MUCA to respond immediately (as described in Section 3.2.3). Upon addressing Amy’s concern about the insurance fee, MUCA’s feedback not only validated her perspective but also offered supplementary insights to elucidate and support her argument. This clarification enhanced Amy’s viewpoint, thereby promoting a more comprehensive grasp of her statement among the other participants in the discussion.
@@
In-Context Chime-in (provide insights for decision-making): In Figure 6-B), it is noteworthy that although MUCA-small was not directly pinged by any participant, it could perceive that the discussion involved selecting between a public park and a learning center for hosting an event. Also, it highlighted other important factors that were not explicitly mentioned by the participants. By doing so, it offered insights and promoted deeper and more comprehensive considerations in the decision-making process.
@@
In-Context Chime-in (provide insights for a stuck conversation): In Figure 6-C), all the participants were indecisive and unsure about the location of the event, leading to a situation where the conversation was essentially stuck, as indicated by the backend log of MUCA-small. Once recognizing this, MUCA provided pros and cons for the two candidate venues and attempted to assist in progressing the conversation.
@@
In-Context Chime-in (answer general questions): In Figure 6-D), Cindy raised a general question which is not for specific participants. By recognizing this, MUCA-small proactively chimed in to provide an answer, which helped the conversation move on to the next phase. In contrast to a single-user system where a chatbot invariably responds to a user’s query, a multi-user chatbot faces a more complex challenge: it has to determine whether to reply to questions when not directly pinged. This suggests that a more subtle decision-making process is required for a multi-user chatbot, as described in Section 3.1.2.
@@
Participation Encouragement: In Figure 6-E), we can see that while Amy, Dennis, and Bob were actively engaged in the discussion, Cindy remained silent for an extended period of time. MUCA-small identified the inactive status of Cindy and hence decided to encourage Cindy to participate based on Cindy’s chatting record. This dialog act tries to enhance Cindy’s engagement, which in turn enables a chat environment with more diverse opinions. We emphasize here that MUCA, by design, is programmed to avoid excessively pinging Cindy or any user who appears reluctant to share their ideas, thereby ensuring a balanced and non-intrusive interaction, as described in Section 3.2.3.
@@
Initiative Summarization: Initiative summarization is generated automatically by MUCA after 12 turns to conclude the existing conversation. In Figure 6-F), we can observe that MUCA-small triggered this function and compiled key points from chatting history. In this example, MUCA not only summarized the diverse opinions of each participant but also offered its own viewpoint, trying to help participants reach an agreement. It empowers participants to quickly catch up with the long and ongoing discussion so that they can contribute their thoughts without the need to review extensive chatting history.
@@
Sub-topic Transition: In Figure 6-G)-1 and G)-2, MUCA-small collected information during the chat, and attempted to shift the topic at an appropriate moment. As mentioned in Section 3.2.3, this dialog act would be triggered in two conditions: (1) As shown in Figure 6-G)-1, currently only a limited number of participants were engaged in a topic that had previously been extensively discussed by the majority, it would attempt to guide the conversation towards the next phase. (2) As shown in Figure 6-G)-2, if MUCA detected that the current topic had been discussed by only a small number of participants, as Amy and Bob had not participated from the beginning, it then attempted to transit the topic to improve user engagement. In both scenarios, MUCA was customized to ping inactive participants and tried to boost their interest according to their chatting history before transiting the topic. Besides, the transition was not abrupt or forceful; it politely inquired whether participants wished to continue or shift the discussion, thus reducing the possibility of any potential disruption. Note that MUCA is designed to guide rather than dictate the flow of dialogue, ensuring that the conversation remains participant-driven.
@@ 4.2.2 Functionality Comparison
As shown in Figure 7, we compared the behaviors of proposed MUCA-small and those of the naive Baseline-small to demonstrate the effectiveness of the design.
@@
Hallucination: As shown in Figure 7-A)-1 and A)-2, Baseline-small can sometimes make up of some assumed information beyond the user-input information (topics, hints, and agenda), such as budget limit, participants’ age, and their genders to answer the questions. These assumptions may lead to user distrust, wrong decisions, and even bias. On the contrary, MUCA-small kindly notified the users that the question was out of scope, and its responses were aligned with the user input information.
@@
We dive deeper into this issue in Figure 7-B)-1 and B)-2. For the unknown budget information, Baseline-small at first provided a non-existing number, which was later questioned and corrected by Bob who may know the real budget. Later, when Cindy tried to figure out who may know the budget, Baseline-small still insisted on this hallucinated information, and even tried to make topic transition. In contrast, for MUCA-small, it inferred that Bob may know the information, as he actively discussed this information. We are highlighting that, unlike single-user chatbots, multi-user chatbots have to process more complicated chat history, participants’ relationship, and user interactions, all of which present challenges for the chatbot to extract useful information and generate hallucination-free response. Addressing the above issues may require dedicated and careful prompting design, even with a powerful LLM as its backend model.
@@
Direct Chatting (passive summarization): Summarization or similar tasks (such as voting) are common requests, especially when users have different opinions in a long conversation. We define this request as passive summarization, in contrast with the Initiative Summarization as mentioned in Section 3.2.3. We compare this functionality in Figure 7-C)-1 and C)-2. For Baseline-small, it failed to understand the query intent from Dennis, which was summarizing the votes from all users. Instead, it summarized opinions, and its summary was inaccurate due to the limited context window by design. For example, it mentioned the "Majority" leaning towards the learning center, where only Dennis voted for this option. On the contrary, MUCA-small, which can go beyond the limit of the window size, correctly summarized the voting results and categorized them by different users.
@@
Figure 8: A detailed comparison from the user study between the Baseline-small and the MUCA-small. The left two and the right two columns represent two topics, indoor courses (Topic-A) and interview process (Topic-B), respectively. Each set of results presents the performance of Baseline-small in the first row, and the performance of MUCA-small in the second row.
@@
Conflict Resolution: It is not unusual that participants have diverse opinions in a multi-user chatting environment, while it is not a very common scenario for a single-user setting. As shown in Figure 7-D)-1, Baseline-small tried to resolve the conflict by instructing users with its own opinion, which can sometimes be misleading and biased. Furthermore, after giving its own opinion, Baseline-small attempted to switch to another topic, which may make users feel annoyed as it interrupted the conversation flow. In Figure 7-D)-2, MUCA-small wrapped up the conflict by different parties, and raised inspiring questions for users to think thoroughly, and resolved conflict when possible.
@@ 4.3 User Study
In this section, we conducted user studies to qualitatively and quantitatively demonstrate the effectiveness of the proposed MUCA-small and MUCA-medium, in comparison with Baseline-small.
@@ 4.3.1 Study Design and Procedure
To investigate MUCA’s effectiveness, we conducted user studies with three participant groups, two small groups (4 people each) and one medium group (8 people), maintaining a 1:1 ratio of females to males. We chose two representative goal-oriented topics for the user studies, as shown in Section 4.1.2. The small group experiments compared Baseline-small with MUCA-small, while the medium group experiment demonstrated the MUCA’s capabilities on more complex chatting scenarios in a larger group.
In small-group experiments, the goal was to evaluate the proposed chatbot (MUCA-small) against a baseline (Baseline-small). Group-A tested Topic A first with Baseline-small then MUCA-small, and Topic B in the reverse order. While Group-B reversed the chatbot order in the experiments to counter the learning effect, where participants might become more familiar with the topic after interacting with the first chatbot. Additionally, MUCA was applied to a medium group using Topic-A, demonstrating its effectiveness in larger group settings.
4.3.2 Qualitative Comparison in Small-size Groups
Statistics from users: Figure  8 presents a qualitative comparison of MUCA-small and the baseline Baseline-small across four aspects:
@@ 4.3.2 Qualitative Comparison in Small-size Groups
Figure 9: General feedback from users for comparison between the MUCA-small and the Baseline-small.
@@
Chime-in Timing: Participants have observed that both chatbots have ever chimed in at the right time during the whole conversation, while MUCA-small performs slightly better, as demonstrated in Figure 8(a). However, Baseline-small also has been observed to frequently chime in at the wrong timing during the chat, unlike MUCA-small, which exhibits this behavior much less frequently. Note that 56.25% (9 out of 16) participants observed that Baseline-small chimes in excessively. This is believed to be a result of its less strategically designed behavior – it always replies every three turns ($N_{exe}=3$) and ignores the "keeping silent" instruction in its prompt, as described in Section 4.1.1. In contrast, MUCA-small has not exhibited this issue as 0% (0 out of 16) participant report it.
However, some participants observe that MUCA-small sometimes chime in insufficiently. This issue might stem from its pre-defined execution interval to maintain latency-efficiency as well as the dialog act of keeping silence, which causes it to reply every three turns at the most if not directly pinged by participants. Although the $N_{exe}$ could be reduced for MUCA-small, this would likely increase both response latency and resource usage, presenting a typical trade-off between resource and performance. We think we may also need more market investigation in the future to decide the best choice of $N_{exe}$ for a production-level design of a multi-user chatbot. One more reason for insufficient chime-in concluded through our observation is as below. In text-based communication systems, unlike speech-based systems, participants might overlook information provided by the chatbot, as evidenced in some chat logs. This oversight can lead to a gap between the information the chatbot has provided and the participants’ awareness. Therefore, sometimes users expect the chatbot to repeat information it has already given.
@@
Chime-in Content: Besides chiming in with good timing, MUCA-small usually produces appropriate content, as shown in Figure 8(b), and inappropriate content is not commonly observed. However, Baseline-small repeats the information it mentions before, asks redundant questions, and tends to generate excessive content. Some of the provided information might be useful but some of them are overwhelmed, which might waste participants more effort to distinguish them.
@@
Participation Encouragement: The interaction feature, i.e., pinging a lurker by a chatbot, should be cautiously designed, including its timing, frequency, and contents, as it may impose negative feelings on some participants due to different personalities, while a good design may improve the efficiency of the chatting and make users more engaged. From this user-friendly perspective, as shown in Figure 8(c), MUCA-small generally has better behaviors with the proposed design, as it improves comfortableness and reduces users’ negative feelings.
@@
As we can see, MUCA-small shows notable successes in most cases as demonstrated in Figure 8(d). From the perspective of the three major user-friendly factors, MUCA-small received significantly higher scores.
  Metrics Indoor Course (Topic-A) Interview Agenda (Topic-B)
B*-sml M*-sml M*-med B*-sml M*-sml
Engt.-Words/Conv. 426.5 531.5 875 531 636.5
Engt.-Words/Utt. 7.23 8.93 8.75 8.85 11.27
Evenness 106.6 $\pm$ 67.6% 132.9 $\pm$ 47.1% 109.4 $\pm$ 56.0% 132.8 $\pm$ 58.0% 159.1 $\pm$ 61.2%
Consensus (%) 50 66.7 66.7 50 100
  Efficiency (%) 50 62.5 68.75 59.38 59.38
Conciseness (%) 31.25 71.88 75 53.13 59.38
Usefulness (%) 43.75 71.88 65.63 65.63 68.75
Overall Rating (%) 37.5 69.44 69.44 52.78 63.89
Table 1: Comparisons among proposed MUCAs and baseline system in terms of quantitative results (upper three rows) and user evaluation scores (bottom four rows). B*-sml, M*-sml, and M*-med represent Baseline-small, MUCA-small, and MUCA-medium, respectively.
Figure 9 provides a direct comparison of participant feedback, preferring MUCA-small’s performance in three key areas. Firstly, MUCA-small can keep the conversation flow without interrupting participants. Secondly, MUCA-small excels as an assistant by summarizing topics, aiding understanding, responding to requests, and enhancing user engagement. Thirdly, MUCA-small effectively provides insights and inspirations for many aspects.
4.3.3 Quantitative Study in Small-size Groups